if jims age is added to sandras age, the sum is 18. if twice jims age is subtracted from sandras age, the difference is 3. how old is each?
Answer by stanbon(75887)

if jims age is added to sandras age, the sum is 18. if twice jims age is subtracted from sandras age, the difference is 3. how old is each?
-----
Equations:
j + s = 18
s - 2j = 3
-------
Substitute for "s" and solve for "j":
18-j - 2j = 3
-3j = -15
j = 5 yrs old (jim's age now)
------
Solve for "s":s = 18-j
s = 13 yrs old (Sandra's age now)
==========================
Cheers,
Stan H.